Find the number of terms in the arithmetic sequence with the given conditions.
a1=-3, d=1/3, S=55
n=

Answer by stanbon(75887)

Find the number of terms in the arithmetic sequence with the given conditions.
a1=-3, d=1/3, S=55
n=
----------------
S(n) = (n/2)(a + L)
----------------
S(n) = (n/2)(a+a+(n-1)d)
------
55 = (n/2)(2a+nd - d)
---
55 = (n/2)(-6 + (1/3)n - (1/3))
----
55 = (n/2)(-19/3) + (1/3)n)
----
55 = (n/2)[(-19+n)/3)
----
55 = (n/6)(-19+n)
----
330 = -19n + n^2
---
n^2 - 19n - 330 = 0
-----
(n+11)(n-30) = 0
---
n = 30
=================
Cheers,
Stan H.
==================